OpenStudy (anonymous):

what is the prime factorization of 135

jimthompson5910 (jim_thompson5910):

First, divide 135 by 5 to get 27 So 135 = 27*5 Now factor 27. You do this by dividing by 3 to get 27/3 = 9 So 27 = 9*3 This means that 135 = 9*3*5 Finally, write 9 as 3*3 to get the answer of 135 = 3*3*3*5